Three atoms of Darstadtium Ds (z=110), were produced. What mass of Ds were produced? (AW of Ds = 269 amu)
1
Expert's answer
2018-09-20T07:25:57-0400
Solution:
1 a.m.u = 1,660 539 040(20) × 10−27kg
m= a.m.u.* Aw = 1.660539040×10-27 * 269= 4,4668499×10-25 kg
Answer: m= 4,4668499×10-25 kg